OpenStudy (anonymous):

how to find the slope intercept form of y+x=-3

OpenStudy (agreene):

y+x=-3 subtract x from both sides y=-3-x do some switching y=-x-3 yay, y=mx+b form :)

OpenStudy (anonymous):

so the answer is (0,3) ?

OpenStudy (agreene):

well, your slope is -1 and your y intercept is -3, if you check for x=0 y=-0-3 y=-3 so (0,-3) is on the line. @ x=5 y=-5-3 y=-8 (5,-8) is on the line as well.

OpenStudy (anonymous):

so i put (-1,-3) on the line ? sorry im so confused